Imagine that the gas shown in the simulation is an
Answer #1

For an ideal gas , PV=nRT

Or,V=nRT/P

Where,R=universal gas constant=0.0821 L atm K-1 mol-1)

T=temperature

(standard temp=273K)

P=pressure (std P=1 atm)

N= number of moles

V1=n1RT/P=(1.00 mol)* RT/P

V2=n2RT/P=(2.00 mol)* RT/P

V1/V2=1/2

2V1=V2

Putting the values of P,R.T,

V1=n1RT/P=(1.00 mol)*(0.0821 L atm K-1 mol-1)*(273K)/1 atm=22.413L

V2=2*V1=2*22.41L=44.82 6L

V2=44.826L